### v2

# SelectTTS: Synthesizing Anyone’s Voice via Discrete Unit-Based Frame Selection

Abstract
Synthesizing the voices of unseen speakers remains a persisting challenge in multi-speaker text-to-speech (TTS). Existing methods model speaker characteristics through speaker conditioning during training, leading to increased model complexity and limiting reproducibility and accessibility. A lower-complexity method would enable speech synthesis research with limited computational and data resources to reach to a wider use. To this end, we propose SelectTTS, a simple and effective alternative. SelectTTS selects appropriate frames from the target speaker and decodes them using frame-level self-supervised learning (SSL) features. We demonstrate that this approach can effectively capture speaker characteristics for unseen speakers and achieves performance comparable to state-of-the-art multi-speaker TTS frameworks on both objective and subjective metrics. By directly selecting frames from the target speaker’s speech, SelectTTS enables generalization to unseen speakers with significantly lower model complexity. Compared to baselines such as XTTS-v2 and VALL-E, SelectTTS achieves better speaker similarity while reducing model parameters by over 8× and training data requirements by 270×.
Index Terms:
Multi-speaker TTS, frame selection, self-supervised learning
I Introduction
Recent text-to-speech approaches [1, 2, 3] have demonstrated that given sufficient data and large enough model capacity, TTS models are capable of producing speech of remarkably high quality and naturalness. While the development of large-scale TTS models [4] brings significant advantages, it also introduces challenges associated with model reproducibility. This issue becomes especially relevant in multi-speaker TTS frameworks. Multi-speaker TTS for unseen speakers is a challenging problem as the objective is multifold - learning text-to-speech semantic prediction while simultaneously capturing the speaker timbre and acoustics[5]. Previously, multi-speaker TTS predominantly relied on speaker labels and embeddings [6, 7] to condition models to learn speaker traits [8, 9]. More recently, zero-shot TTS frameworks have advanced this area by leveraging neural codec language modeling [2, 10] combined with acoustic prompting for speaker timbre control.
Both the embedding and language modeling approach rely on speaker conditioning, either in the form of speaker embeddings or acoustic prompts, requiring the network to learn the conditional modeling of speaker timbre. While leading to state-of-the-art (SOTA) performance [11, 12, 13], it comes at the cost of increased model complexity, capacity, and data requirements [11, 14], limiting reproducibility and further development. In particular, these models are computationally expensive to train and deploy, and their reliance on large-scale computational and data resources makes them impractical for many research and application settings. To address these limitations, we propose SelectTTS, a simple and effective alternative to conditional learning of speaker traits where speaker modeling is performed by non-parametric frame selection of SSL features. This approach enables training more parameter-efficient and less complex models without extensive data requirements, while still achieving speaker timbre reproduction comparable to larger models.
This is made possible by SSL speech models [15, 16], which have demonstrated strong capabilities in capturing linguistic, speaker, and prosodic information [17]. The masked prediction objectives used in these models [18] encourage the learning of rich semantic representations while preserving speaker identity. A recent work, kNN-VC [19], has leveraged this property by performing unit selection-based voice conversion, replacing each frame-level feature in the source utterance with its nearest neighbors in the reference speech. This frame-by-frame selection of SSL features enables the synthesis of voices with high similarity to the target speaker.
Speech Samples: https://kodhandarama.github.io/selectTTSdemo/Codes and pre-trained models will be released upon acceptance.
Figure 1: *
(a) Semantic unit tokenizers
Figure 2: *
(b) SelectTTS inference
Figure 3: Proposed SelectTTS framework with the frame-selection method. In the frame selection, frames $z_{1}$,$z_{2}$,$z_{3}$,$z_{4}$ are chosen through sub-sequence matching and frames $z_{7}$, $z_{9}$,$z_{6}$ and $z_{10}$ are chosen via inverse k-means sampling.
With SelectTTS, we introduce a new paradigm of frame selection-based multi-speaker TTS that directly utilizes frames from unseen target speakers to clone their voice. Instead of conditioning the input text with speaker embeddings or using speech tokens as prompts to model speaker timbre, SelectTTS splits the task into two stages. In the first stage, frame-level semantic units are predicted from text to model the speech content. In the second stage, intermediate representations enriched with speaker and acoustic information are obtained by selecting frame-level features from the reference speech based on the predicted semantic units. Combining frame selection algorithms with rich SSL features achieves SOTA speaker similarity, as it directly leverages frames from the target speaker. This multi-stage, non-parametric framework reduces model complexity and offers greater customization compared to larger baseline systems.
The main contributions of this work are as follows: 1) We propose a multi-speaker TTS strategy that completely separates and simplifies the tasks of semantic prediction and speaker modeling in TTS, making the overall framework easily reproducible and open to further development; 2) We introduce novel frame selection algorithms, sub-sequence matching and inverse k-means sampling, that directly select frames from the target speaker to accurately reproduce speaker timbre; 3) We utilize the discrete SSL features for text-to-semantic prediction and frame selection; the continuous SSL features for vocoding, showing the benefits of leveraging both feature spaces.
II Related work
Learning speaker characteristics has been a primary focus of multi-speaker TTS research. In the deep learning era, traditional approaches have primarily relied on conditioning neural networks with speaker representations [20, 21, 3, 7, 22]. More recently, inspired by large language models (LLMs), in-context learning with large codec-based models, such as VALL-E[2], VoiceCraft[11], Voicebox[1] have been explored. These models leverage bi-directional context for speech-infilling tasks using speaker information in the form of acoustic prompts, achieving SOTA speaker similarity with the cost of training in very large-scale data. Methods such as SPEAR-TTS[23], UniCATS [24] reduce the data requirement compared to those models while using acoustic prompting, yet, they remain to be complex. Additionally, methods such as UnitSpeech[25], HierSpeech[26] perform speaker adaptation by fine-tuning. In this work, we explore an alternative direction and show that a simple, non-parametric frame selection method can achieve SOTA speaker similarity without the complexity of large-scale models.
III SelectTTS
SelectTTS consists of two training stages and one offline intermediate stage. In the first training stage, we train a text-to-semantic-unit prediction model. Using the predicted semantic units, we perform frame selection by identifying corresponding frames from the target speaker. This offline frame selection step enables the recovery of continuous SSL features from their discrete counterparts, as illustrated in Fig. 1(b). Finally, in the second training stage, we train a vocoder to map the selected frame-level continuous SSL features to the speech waveform. In the following subsections, we describe each stage of the framework in detail.
III-A Semantic unit tokenizers
SelectTTS relies on its frame selection pipeline to choose the correct frames from the reference speech. To enable effective matching, frame selection is performed in a discrete semantic-unit space.
III-A1 Speech-unit-tokenizer
The speech-unit tokenizer converts speech into frame-level sequences of semantic units. Continuous SSL features are extracted from the speech at the frame level using a pre-trained SSL model. These features are then converted into discrete semantic units through k-means clustering, as in Fig. 1(a). We denote this transformation as $z$ = SpeechUnitTokenizer($Z$), where $z$ represents the discrete unit sequence and $Z$ is the continuous SSL feature sequence.
III-A2 Text-unit-tokenizer
We train a non-autoregressive model to predict frame-level discrete semantic units directly from text. For this task, we adopt an architecture based on FastSpeech2 [27], consisting of a text encoder, a duration prediction module, and a unit decoder. The model is trained on a parallel dataset containing text and corresponding discrete speech units, where the speech units are obtained using the SpeechUnitTokenizer transform. An external alignment tool [28] is used to obtain phoneme durations, which serve as supervision for training the duration predictor.
During inference, the TextUnitTokenizer consumes non-aligned phoneme sequences and predicts frame-level sequences of discrete semantic units. The duration model predicts phoneme durations and upsamples the phoneme sequence accordingly to match the frame-level granularity. We emphasize that predicting discrete units instead of continuous features simplifies the modeling task and reduces overall system complexity. Moreover, operating in the discrete space enables more effective frame selection, allowing sequences of frames to be selected together, thereby capturing intra-frame dependencies that would be difficult to handle with continuous features.
III-B Frame selection algorithms
We propose a novel frame selection pipeline that leverages two algorithms in sequence: sub-sequence matching followed by inverse k-means sampling. The pipeline selects an appropriate frame from the reference speech corresponding to each predicted unit produced by the TextUnitTokenizer.
III-B1 Sub-sequence matching
We formulate the sub-sequence matching problem as follows: Let $\hat{z}$ = TextUnitTokenizer(text) represent a sequence of predicted semantic units obtained from the TextUnitTokenizer. Similarly, let $z_{\text{ref}}$ = SpeechUnitTokenizer(speech) denote the semantic unit sequence derived from the reference speech of the target speaker using the speech-unit tokenizer. In our sub-sequence matching algorithm, we find sub-sequences of the predicted sequence $\hat{z}$ that is present in the reference unit sequence $z_{\text{ref}}$. Once we find these sub-sequences, we replace the sub-sequence-matched discrete unit sequence with the corresponding continuous SSL features from the reference speech as in Fig. 1(b). We start from the longest sub-sequence length and iteratively replace the discrete units till there are no more matches. In the implementation, we begin by searching for matches with a maximum length of 10 to maintain computational efficiency, with a minimum length of 2. The intuition of sub-sequence matching is that by choosing chunks of speech segments at a time, we hope to get the most accurate match in the form of real speech segments instead of relying on only frame-level selection, to reduce artifacts and improve segment-level prosody.
III-B2 Inverse k-means sampling
For frames without direct sub-sequence matches, we apply a strategy called inverse $k$-means sampling. In this approach, we identify frames from the reference speech that belong to the same discrete unit cluster as the predicted semantic unit. This effectively performs the reverse of the $k$-means clustering process by retrieving continuous SSL features from the discretized semantic space. If the predicted unit cluster is not represented in the reference speech - typically due to the limited reference duration - we instead select frames from the nearest nonempty cluster. Once the appropriate cluster is determined, the corresponding continuous SSL feature is obtained either by randomly sampling a frame from the target cluster or by averaging all SSL features associated with that cluster in the reference speech.
III-C Vocoder
The vocoder converts continuous WavLM features into audio waveforms. It can be trained with ground-truth frame-level feature and speech waveform pairs, independently from frame selection algorithms. However, this causes a mismatch between the inputs during training and inference. To overcome this mismatch, we perform frame selection with the ground truth audio before training the vocoder. Given an input utterance in a set of utterances from a speaker, we choose the reference speech to be every other utterance from the same speaker. We extract ground truth WavLM units using SpeechUnitTokenizer from the input utterance and perform frame selection described in section III-B. The selected frame-level SSL features are then used as input to the vocoder. The target for the vocoder is the ground-truth audio. In preliminary experiments, we observe that fine-tuning the vocoder using subsequence-matched features significantly reduces synthesis artifacts and improves audio quality.
III-D SelectTTS vs kNN-VC
SelectTTS is inspired by the retrieval-based VC method kNN-VC [19], but it introduces a fundamentally different task and framework. Unlike SSL feature-based VC, where SSL features can be directly extracted from input speech, SelectTTS predicts semantic representations from text and performs frame selection from reference speech to synthesize new utterances. This design bridges text and speaker-specific SSL features, enabling efficient, low-complexity TTS for unseen speakers.
IV Experimental setup
IV-A SelectTTS implementation
The choice of SSL feature layer is critical in our framework. We initially experimented with layers 22 and 23 of WavLM-Large [16], which are effective for phoneme recognition due to their rich linguistic content. However, features from these higher layers degraded speaker identity and prosody reproduction, both essential for multi-speaker TTS [29]. Based on this observation, we adopt layer 6 of WavLM-Large, following kNN-VC, as it better preserves speaker-specific information.
IV-A1 Speech-unit-tokenizer
The WavLM-Large encoder produces continuous feature vectors for every 20ms of 16kHz of audio. We discretize these features using $k$-means clustering, as described in Section III-A1. For our proposed method, we use 2000 cluster centers. This choice was made after initial experiments with 100 and 500 clusters, where we observed that more fine-grained quantization led to improved intelligibility of the synthesized speech.
IV-A2 Text-unit-tokenizer
We modify the FastSpeech2 architecture [27] to design a text-to-semantic-unit predictor, replacing Mel spectrogram prediction with semantic unit prediction. The ground truth units are derived by applying the SpeechTokenizer on the ground truth audio. The network is optimized using the Adam optimizer with a learning rate of $45e-$ 4 and a batch size corresponding to 10,000 phonemes. Training is performed on an NVIDIA RTX 3090 GPU, where the network converges within approximately 20 minutes, or after around 10,000 steps. Our implementation builds upon the publicly available SpeechLM framework [30].
3) Vocoder: We use a HiFi-GAN V1 architecture from [31]. We start with the pre-trained vocoder model from the kNN-VC implementation [^1] https://github.com/bshall/knn-vc. We fine-tune two vocoders with frames selected by proposed algorithms: one where frame selection is based solely on inverse k-means sampling, and another where frame selection incorporates both inverse k-means sampling and sub-sequence matching.
IV-B Baselines
We use three SOTA multi-speaker TTS baselines to compare the performance of our proposed framework: the official releases of YourTTS [^2] https://github.com/Edresson/YourTTS [7] and XTTS-v2333https://huggingface.co/coqui/XTTS-v2 [3] with default parameters. We use Amphion’s VALL-E  [^4] https://github.com/open-mmlab/Amphion/tree/main/egs/tts/VALLE_V2 [2] implementation.
IV-C Training and test data
We utilize LibriSpeech train-clean-100 dataset [32] for training both the text-to-semantic-unit tokenizer and the HiFi-GAN vocoder. This dataset comprises 100 hours of text-audio pairs, sampled at 16kHz, and features 251 speakers. We use the LibriSpeech dev-clean subset for validation and LibriTTS-R test-clean [33] for testing. For the LibriTTS-R test-clean set, we select speakers who were not seen during training and have at least 20 utterances, each containing between 3 and 30 words. This results in 31 eligible speakers and a total of 620 utterances. For the VCTK [34] subset which is unseen during training, we sampled 60 speakers where each speaker utters 10 randomly selected sentences, resulting in a total of 600 utterances. We use 5 minutes of reference audio for all variants of SelectTTS, as well as for YourTTS and XTTS-v2, across all evaluation settings. For VALL-E, due to computational constraints, the reference speech was limited to 10-15 seconds.
TABLE I: Objective evaluation results (LibriTTS-R test-clean)
Method WER(%$\downarrow$ SECS $\uparrow$ UTMOS $\uparrow$
Ground Truth 3.55 68.12 4.22
VALL-E 4.72 58.24 4.08
XTTS-v2 4.23 60.26 4.16
YourTTS 9.66 50.83 3.61
SelectTTS (only inv k-means (rand)) 7.19 62.84 3.46
SelectTTS (only inv k-means (avg)) 6.49 64.74 3.89
SelectTTS (inv k-means (rand) + sub-match) 7.31 61.57 3.99
SelectTTS (inv k-means (avg) + sub-match) 6.67 61.59 4.13
TABLE II: Objective evaluation results (VCTK)
Method WER(%$\downarrow$ SECS $\uparrow$ UTMOS $\uparrow$
Ground Truth 4.39 69.77 4.04
VALL-E 8.53 46.70 4.07
XTTS-v2 4.89 54.12 4.00
YourTTS 9.90 36.83 3.53
SelectTTS (only inv k-means (avg)) 7.18 60.61 3.69
SelectTTS (inv k-means (avg) + sub-match) 7.01 54.13 3.92
TABLE III: Analysis of objective performance with varying reference speech lengths
WER(%$\downarrow$ SECS $\uparrow$
Method 3min 1min 30s 3min 1min 30s
SelectTTS w/o sub-match 6.55 8.86 13.35 63.85 62.17 58.62
SelectTTS w sub-match 6.99 8.87 11.73 60.96 59.96 57.34
V Results and discussion
We evaluate four variants of the proposed SelectTTS framework. The first variant, inv-kmeans (rand), employs inverse $k$-means sampling by randomly selecting a frame from the predicted cluster. The second variant, inv-kmeans (avg), also uses inverse $k$-means sampling but decodes using the average of all frames within the predicted cluster. The third and fourth variants build on the first two by incorporating sub-sequence matching.
V-A Objective evaluation
Table I presents objective evaluation results on the LibriTTS-R test set, measuring intelligibility via WER using the Wav2Vec 2.0 Large ASR model [^5] https://huggingface.co/facebook/wav2vec2-large-960h-lv60-self, and naturalness via UTMOS [35]. We observe that sub-sequence matching improves both metrics. SelectTTS outperforms YourTTS in intelligibility and naturalness, and approaches the performance of larger models like VALL-E and XTTS-v2. SelectTTS also achieves the highest speaker similarity across all models, as measured by SECS using ECAPA-TDNN embeddings [36]. Results on the VCTK test set (Table II) show a similar trend: the proposed method attains the best speaker similarity with competitive WER and UTMOS scores, demonstrating robustness across datasets. Table III analyzes the impact of reference speech duration. Although performance slightly drops with shorter references, SelectTTS remains competitive even with only 30 seconds of reference audio. As shown in Table IV, SelectTTS requires 8× fewer parameters than XTTS-v2 and 10× fewer than VALL-E, with 270× and 450× less training data, respectively, while maintaining a comparable Real-Time Factor (RTF).
TABLE IV: Analysis of model complexity
Method #Params Train data (hours) RTF $\downarrow$
VALL-E 594M 45k 1.325
XTTS-v2 466M 27k 0.186
YourTTS 86M 474 0.120
SelectTTS w/o sub-match (inv k-means (avg)) 57M 100 0.290
SelectTTS w sub-match (inv k-means (avg)) 57M 100 0.334
TABLE V: Subjective evaluations (95% confidence interval)
[…]
Ground Truth 4.20 $\pm$ 0.09 -
VALL-E 4.11 $\pm$ 0.08 3.62 $\pm$ 0.09
XTTS-v2 4.10 $\pm$ 0.08 3.64 $\pm$ 0.10
YourTTS 3.15 $\pm$ 0.10 3.22 $\pm$ 0.09
SelectTTS w/o sub-match (inv k-means (avg)) 3.89 $\pm$ 0.08 3.71 $\pm$ 0.09
SelectTTS w sub-match (inv k-means (avg)) 3.97 $\pm$ 0.08 3.72 $\pm$ 0.09
V-B Subjective evaluation
We conduct listening experiments with 21 participants to evaluate the naturalness and speaker similarity of the synthesized speech. A total of 120 utterances from 6 unseen speakers (3 male, 3 female) from the LibriTTS-R test set are used for the evaluation. In the Mean Opinion Score (MOS) test [7], participants rate the naturalness of speech samples on a 5-point scale, evaluating 12 utterances per method. As shown in Table V, SelectTTS achieves a high MOS score close to 4, comparable to strong baselines. Sub-sequence matching further improves the perceived naturalness. We also conduct a Speaker Mean Opinion Score (SMOS) evaluation [7], where participants are presented with a ground-truth recording and a synthesized utterance and asked to rate speaker similarity on a 5-point scale. SelectTTS achieves strong SMOS results, demonstrating speaker resemblance on par with or better than larger baseline models.
VI Conclusion
In this paper, we propose SelectTTS, a multi-speaker TTS framework with lower model complexity that synthesizes high-quality speech closely resembling the target speaker’s voice by directly utilizing frames from unseen speakers. We demonstrate that combining semantic unit-based frame selection with vocoding over SSL feature sequences provides a simple yet effective approach for modeling unseen speakers, achieving state-of-the-art speaker similarity. SelectTTS significantly reduces model complexity and data requirements, opening possibilities for extending state-of-the-art multi-speaker TTS to low-resource settings and languages.